OpenStudy (anonymous):

solve for x: 1/3x+1/2=-2/3x+2/5

OpenStudy (anonymous):

1) multiply both sides by 3x:\[3x(\frac{ 1 }{ 3x }+\frac{ 1 }{ 2 })=(\frac{ -2 }{ 3x }+\frac{ 2 }{ 5 })\] to get \[\frac{ 3x }{ 2 }+1=\frac{ 6x }{ 5 }-2\] Getting a common denominator: \[\frac{ 5 }{ 5 }(\frac{ 3x }{ 2 }+1)=\frac{ 2 }{ 2 }(\frac{ 6x }{ 5 }-2)\] This is allowed since 5/5 and 2/2 are actually just 1, so they don't affect the equation at all. So now we get \[\frac{ 15x }{ 10 }+1=\frac{ 12x }{ 10 }-2\] Reducing:\[\frac{ 3x }{ 10 }=-3\] Giving us: \[3x=-30\] and \[x=-10\]

OpenStudy (anonymous):

find the common denominator of all four denominators (3,2,3,5) which is 30. Multiply the equation by 30 to get rid of the fractions... (30)1/3x + (30)1/2= (30)-2/3x + (30)2/5 10x + 15 = -20x + 12 (now get x's on one side, everything else on the other ) 10x + 20x = 12 - 15 30x = - 3 (divide by 30 to get x by itself) 30x/30 = - 3/30 x = - 1/10 check... 10x + 15 = - 20x + 12 10(-1/10) + 15 = - 20(-1/10) + 12 -1 + 15 = 2 + 12 14 = 14 (correct)